Is it more profitable to receive $2000 now or $3400 in 9 years, if the money can earn 6% interest compounded quarterly?
1
Expert's answer
2020-06-28T16:50:35-0400

"Present \, Value \, PV =FV (1+i)^{-n}" ,

i= 0.06/4= 0.015,

PV=3400(1+0.15)-36 =$ 1,989.31

So the value of $3400 received in 9 years is less than $2000 received today. The better option for selection is to receive $2000 today.